# AI Foundation Models in Remote Sensing: A Survey

### v1

@@ Abstract
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing data collection, processing, and analysis. Traditionally reliant on manual interpretation and task-specific models, remote sensing has been significantly enhanced by the advent of foundation models—large-scale, pre-trained AI models capable of performing a wide array of tasks with unprecedented accuracy and efficiency. This paper provides a comprehensive survey of foundation models in the remote sensing domain, covering models released between June 2021 and June 2024. We categorize these models based on their applications in computer vision and domain-specific tasks, offering insights into their architectures, pre-training datasets, and methodologies. Through detailed performance comparisons, we highlight emerging trends and the significant advancements achieved by those foundation models. Additionally, we discuss the technical challenges, practical implications, and future research directions, addressing the need for high-quality data, computational resources, and improved model generalization. Our research also finds that pre-training methods, particularly self-supervised learning techniques like contrastive learning and masked autoencoders, significantly enhance the performance and robustness of foundation models in remote sensing tasks such as scene classification, object detection, and other applications. This survey aims to serve as a resource for researchers and practitioners by providing a panorama of advances and promising pathways for continued development and application of foundation models in remote sensing.
@@ I Introduction
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing how data is collected, processed, and analyzed. Traditionally, remote sensing projects relied heavily on manual interpretation and task-specific models that required extensive labeled datasets and significant computational resources. However, the advent of AI and deep learning (DL) has ushered in a new era in which large-scale, pre-trained models, known as foundation models, are capable of performing a wide array of tasks with unprecedented accuracy and efficiency. These advancements have not only enhanced the capabilities of remote sensing but have also opened new avenues for its application across various domains.
In recent years, numerous foundation models have emerged, demonstrating remarkable performance in handling diverse remote sensing tasks. These models have shown the potential to significantly improve performance on multiple downstream tasks such as scene classification, semantic segmentation, object detection, and more. By leveraging vast amounts of pre-training data and sophisticated architectures, these foundation models have set new benchmarks in the field, making them indispensable tools for researchers and engineers alike.
This paper aims to provide a comprehensive survey of foundation models in the remote sensing domain, covering foundation models released between June 2021 and June 2024. In figure 1, 51 vision models are listed in chronological order. To facilitate easy navigation and utility for researchers, we have categorized these models based on their application in computer vision tasks and domain-specific tasks. This categorization allows for a clearer understanding of which models are suited for certain purposes, whether they involve general image-based challenges or more specialized applications such as environmental monitoring, land cover and land use, archaeological prospection, disaster management, or other domains.
@@
    An exhaustive review of current state of foundation models proposed in the field of remote sensing, starting from the background and methodologies of foundation models to specific applications across different domains and tasks in a hierarchical and structured manner.
@@
    Categorization and analysis of the models based on their application in both computer vision tasks (table I) and domain-specific tasks (table 5). We discuss the architecture, pre-training datasets, pre-training methods, and performances of each model.
@@
Figure 2: Examples of data types used in those foundation models and downstream tasks that can be done by foundation models. Data: (1) Panchromatic[2], (3) True Color, (4) SAR[68], (5) Hyperspacial[2], (6) Multispectral[2]. Downstream tasks: (1) Segmentation, (2) Object Detection, (3) Classification[10], (4) Change Detection[53].
 [^1] True Color, Segmentation, and Object detection images © MAXAR 2024, provided through the NextView License Agreement.
@@ II-A Remote Sensing
Remote sensing (RS) refers to the process of acquiring information about objects or areas from a distance, typically using satellite or airborne sensors. These technologies and techniques serve vital roles in diverse fields, enabling the collection of data over land geographic areas without physical contact. Applications of remote sensing include earth observation, digital archaeology, urban planning and development, and disaster management. The field of remote sensing has been developed rapidly since the mid 20th century. Initially, remote sensing relied on analog photographic techniques via aerial and satellite platforms, which provided limited spectral and spatial resolution. The launch of early Earth observation satellites, such as Landsat program commenced in 1967 [78], marked a significant advancement, enabling consistent and wide-ranging data collection for environmental monitoring.
Modern remote sensing employs a variety of sensors, including optical, thermal, and radar, each suited for specific types of data collection. Optical sensors capture a wide variety of spectral bands, including visible and near-infrared light, allowing for detailed imaging of land cover and vegetation health. Thermal sensors detect heat emitted or reflected from the Earth’s surface, useful for monitoring volcanic activity, forest fires, and climate change monitoring. Radar sensors can penetrate clouds and vegetation, providing critical information in all-weather conditions and for applications such as soil moisture estimation and urban infrastructure mapping [50, 12].
Year-Month Architecture Model Name Scene Classification Semantic Segmentation Object Detection Change Detection
2021 Jun ResNet-50 CMC-RSSR [59] $\checkmark$
@@
2021 Dec ResNet-34 MATTER [1] $\checkmark$ $\checkmark$ $\checkmark$
2022 Mar ResNet-50 GASSL [3] $\checkmark$ $\checkmark$ $\checkmark$
@@
2022 Jun ViT-S/8 DINO-MM [73] $\checkmark$
2022 Jun Swin Transformer Scheibenreif, et al. [56] $\checkmark$ $\checkmark$
2022 Jul ViT/Swin Transformer RingMo [61] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Aug ResNet-50 GeCO [40] $\checkmark$ $\checkmark$ $\checkmark$
2022 Sep BYOL RS-BYOL [32] $\checkmark$ $\checkmark$
2022 Nov ViT-B CSPT [87] $\checkmark$ $\checkmark$
2022 Nov ViT RVSA [71] $\checkmark$ $\checkmark$ $\checkmark$
2023 Jan MAE-based Framework SatMAE [11] $\checkmark$ $\checkmark$
2023 Apr TOV TOV [63] $\checkmark$ $\checkmark$ $\checkmark$
@@
2023 Jun CACo CACo [47] $\checkmark$ $\checkmark$ $\checkmark$
2023 Jun ResNet-18 IaI-SimCLR [54] $\checkmark$
2023 Aug Teacher-Student GFM [48] $\checkmark$ $\checkmark$ $\checkmark$
@@
2023 Sep Multi-Branch RingMo-Sense [85] $\checkmark$
2023 Sep ViT Scale-MAE [55] $\checkmark$ $\checkmark$
2023 Sep CNN-Transformer RingMo-lite [76] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Sep Multimodel SSL DeCUR [72] $\checkmark$ $\checkmark$
2023 Oct MSFE+MMFH Feng et al. [21] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Oct ViT FG-MAE [75] $\checkmark$ $\checkmark$
2023 Nov ViT Prithvi [33] $\checkmark$
2023 Nov Multimodal Encoder CROMA [22] $\checkmark$ $\checkmark$
2023 Dec ViT USat [31] $\checkmark$
2024 Jan ViT-B Cross-Scale MAE [62] $\checkmark$ $\checkmark$
@@
EarthPT [58] $\checkmark$
2024 Jan Teacher-Student Network GeRSP [29] $\checkmark$ $\checkmark$ $\checkmark$
2024 Jan Dual-Branch SwiMDiff [65] $\checkmark$ $\checkmark$
2024 Jan Generative ConvNet SMLFR [17] $\checkmark$ $\checkmark$
2024 Feb 3D GPT SpectralGPT [28] $\checkmark$ $\checkmark$ $\checkmark$
2024 Feb MAE-based Framework Presto [66] $\checkmark$
2024 Mar SatMAE SatMAE++ [52] $\checkmark$
@@
SAR-JEPA [41] $\checkmark$
2024 Mar ViT FoMo-Bench [5] $\checkmark$ $\checkmark$ $\checkmark$
@@
SkySense [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Mar Multi-Modules UPetu [18] $\checkmark$ $\checkmark$ $\checkmark$
2024 Apr Swim Transformer msGFM [25] $\checkmark$ $\checkmark$
2024 Apr DINO DINO-MC [77] $\checkmark$ $\checkmark$
2024 May OFA-Net OFA-Net [84] $\checkmark$ $\checkmark$
@@
2024 May ViT BFM [7] $\checkmark$ $\checkmark$
2024 May MP-MAE MMEarth [51] $\checkmark$ $\checkmark$
2024 May ViT CtxMIM [86] $\checkmark$ $\checkmark$ $\checkmark$
2024 May HiViT SARATR-X [37] $\checkmark$ $\checkmark$
2024 Jun Dynamic OFA DOFA [83] $\checkmark$
TABLE I: Summery of computer-vision downstream tasks tested by each models.
In recent years, remote sensing has found applications in many fields. In environmental monitoring, it is used to track deforestation, to monitor air and water quality, and to assess the impacts of climate change [23]. In agriculture, remote sensing helps in crop health monitoring, yield estimation, and efficient resource management [50]. Urban planning and development benefit from remote sensing through the monitoring of urban sprawl, infrastructure development, and land-use planning [12]. Furthermore, in disaster management, remote sensing is crucial for assessing the damage caused by natural disasters, aiding in the planning and execution of relief operations [23].
@@ II-B Foundation Models for Remote Sensing
The evolution of FMs has been driven by advancements in deep learning and the availability of large datasets. Initially, convolutional neural networks (CNNs) like ResNet [26] paved the way for improved image recognition and classification tasks [45]. The introduction of transformers, which use self-attention mechanisms to model long-range dependencies, has further advanced the capabilities of FMs in handling large-scale image data [11].
FMs in remote sensing are characterized by their ability to leverage vast amounts of unlabeled data through SSL techniques, allowing them to learn robust representations without the need for extensive labeled datasets [88]. Key SSL methods include contrastive learning, which involves learning representations by comparing different augmented views of the same data point, and predictive coding, which trains models to predict missing parts of the input data from the observed parts [55].
@@
III Methodologies
In recent years, the development of foundation models (FMs) for remote sensing has leveraged various sophisticated methodologies. These include self-supervised learning (SSL), transformers and vision transformers (ViT), and residual neural networks (ResNet). These methodologies have significantly enhanced the capabilities of FMs by enabling them to learn from large amounts of data without extensive human supervision, handle complex data structures, and improve feature extraction and representation. This section reviews the mechanisms and contributions of these methodologies to the field of remote sensing.
III-A Self-Supervised Learning (SSL) Training Strategies for Remote Sensing FMs
Self-supervised learning (SSL) plays a crucial role in the pre-training stage of a foundation model. Through SSL, a model learns to predict part of the input from other parts of the input data, reducing the reliance on large labeled datasets. SSL is particularly valuable in remote sensing due to the scarcity of labeled data. Models pre-trained with SSL can effectively capture patterns and features from vast amounts of unlabeled remote sensing data, making them highly efficient for downstream tasks. Figure 2 presents a general pipeline of self-supervised learning.

### v2

# Vision Foundation Models in Remote Sensing: A Survey

@@ Abstract
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing data collection, processing, and analysis. Traditionally reliant on manual interpretation and task-specific models, remote sensing research has been significantly enhanced by the advent of foundation models—large-scale, pre-trained AI models capable of performing a wide array of tasks with unprecedented accuracy and efficiency. This paper provides a comprehensive survey of foundation models in the remote sensing domain. We categorize these models based on their architectures, pre-training datasets, and methodologies. Through detailed performance comparisons, we highlight emerging trends and the significant advancements achieved by those foundation models. Additionally, we discuss technical challenges, practical implications, and future research directions, addressing the need for high-quality data, computational resources, and improved model generalization. Our research also finds that pre-training methods, particularly self-supervised learning techniques like contrastive learning and masked autoencoders, remarkably enhance the performance and robustness of foundation models. This survey aims to serve as a resource for researchers and practitioners by providing a panorama of advances and promising pathways for continued development and application of foundation models in remote sensing.
@@ I Introduction
Artificial Intelligence (AI) technologies have profoundly transformed the field of remote sensing, revolutionizing how data is collected, processed, and analyzed. Traditionally, remote sensing projects relied heavily on manual interpretation and task-specific models that required extensive labeled datasets and significant computational resources. However, the advent of AI and deep learning (DL) has ushered in a new era in which large-scale, pre-trained models, known as foundation models, are capable of performing a wide array of tasks with unprecedented accuracy and efficiency. These advancements have not only enhanced the potential applications of remote sensing but have also opened new avenues for its usage across various domains.
In recent years, numerous vision foundation models have emerged, demonstrating remarkable performance in handling diverse remote sensing tasks. These models have shown the potential to significantly improve performance on multiple downstream tasks such as scene classification, semantic segmentation, object detection, and more. By leveraging vast amounts of pre-training data and sophisticated architectures, these foundation models have set new benchmarks in the field, making them indispensable tools for researchers and engineers alike.
This paper aims to provide a comprehensive survey of vision foundation models in the remote sensing domain ad rem, and is limited to foundation models released between June 2021 and June 2024. This timeframe marks a surge in the development of modern foundation models, including vision transformers and advanced self-supervised learning techniques. Although early models like Tile2Vec [47] and others laid the groundwork for representation learning in remote sensing, they were typically limited in scale and generalization capabilities. Furthermore, numerous review papers have already provided comprehensive overviews of these pre-2021 models. Our review, therefore, focuses on recent developments to highlight the unique contributions and innovations that have emerged in the past few years.
In figure 1, 58 vision foundation models are listed in chronological order. To facilitate navigation and enhance utility for researchers, we categorized existing models based on their perception levels (e.g., image-level, region-level, pixel-level). This organization helps clarify which models have been tested for general image-based challenges or specialized applications such as environmental monitoring, land cover mapping, archaeological exploration, disaster management, and more. It is essential to distinguish between applications that models have been explicitly tested on and those for which they could potentially be effective for. In this review, the fact that a model has not been tested on a particular application does not mean it won’t perform well. Foundation models, especially convolutional neural network (CNN) backbones like residual networks (ResNet)[36] and vision transformers (ViT) [25], may still be suitable for various downstream tasks, even if prior work has not yet demonstrated this.
@@
    An exhaustive review of current state of vision foundation models proposed in the field of remote sensing, starting from the background and methodologies of these models to specific applications across different domains and tasks in a hierarchical and structured manner.
@@
    Categorization and analysis of the models based on their application in both image analysis (table I) and practical applications (table III). We discuss the architecture, pre-training datasets, pre-training methods, and performance of each model.
@@
Figure 2: Examples of data types used in those foundation models and downstream tasks that can be done by foundation models. Data: (1) Panchromatic[4], (2) True Color, (3) SAR[94], (4) Hyperspectral[4], (5) Multispectral[4]. Downstream tasks: (1) Segmentation, (2) Object Detection, (3) Classification[15], (4) Change Detection[76].  [^2] True Color, Segmentation, and Object detection images © MAXAR 2024, provided through the NextView License Agreement.
@@ II-A Remote Sensing
Remote sensing (RS) refers to the process of acquiring information about objects or areas from a distance, typically using satellite or airborne sensors. These technologies and techniques serve vital roles in diverse fields, enabling the collection of data over geographic areas without physical contact. Applications of remote sensing include earth observation, digital archaeology, urban planning and development, and disaster management. The field of remote sensing has developed rapidly since the mid 20th century. Initially, remote sensing predominately consisted of analog photographic techniques via aerial and satellite platforms, which provided limited spectral and spatial resolution. The launch of early Earth observation satellites, such as Landsat program commenced in 1967 [112], marked a significant advancement, enabling consistent and wide-ranging data collection for environmental monitoring.
Modern remote sensing employs a variety of sensors suited for specific types of data collection, including optical, thermal, and radar.. Optical sensors capture a wide variety of spectral bands, including visible and near-infrared light, allowing for detailed imaging of land cover and vegetation health. Thermal sensors detect heat emitted or reflected from the Earth’s surface, useful for monitoring volcanic activity, forest fires, and climate change monitoring. Radar sensors can penetrate clouds and vegetation, providing critical information in all-weather conditions and for applications such as soil moisture estimation and urban infrastructure mapping [71, 17].
Year-Month Architecture Model Name Image-Level Pixel-Level Region-Level Spatial-Temporal Contrastive Learning Predictive Coding
2021 Jun ResNet-50 CMC-RSSR [84] $\checkmark$ $\checkmark$
@@
2021 Dec ResNet-34 MATTER [2] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Mar ResNet-50 GASSL [6] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
2022 Jun ViT-S/8 DINO-MM [105] $\checkmark$ $\checkmark$
2022 Jun Swin Transformer Scheibenreif, et al. [79] $\checkmark$ $\checkmark$ $\checkmark$
2022 Jul ViT/Swin Transformer RingMo [87] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Aug ResNet-50 GeCO [57] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2022 Sep BYOL RS-BYOL [45] $\checkmark$ $\checkmark$ $\checkmark$
2022 Nov ViT-B CSPT [124] $\checkmark$ $\checkmark$ $\checkmark$
2022 Nov ViT RVSA [100] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Jan MAE-based Framework SatMAE [16] $\checkmark$ $\checkmark$ $\checkmark$
2023 Apr TOV TOV [89] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
2023 Jun CACo CACo [67] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Jun ResNet-18 IaI-SimCLR [77] $\checkmark$ $\checkmark$
2023 Jun ResNet SSL4EO-L [83] $\checkmark$ $\checkmark$
2023 Aug Teacher-Student GFM [69] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
2023 Sep Multi-Branch RingMo-Sense [119] $\checkmark$ $\checkmark$
2023 Sep ViT Scale-MAE [78] $\checkmark$ $\checkmark$ $\checkmark$
2023 Sep CNN-Transformer RingMo-lite [109] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Sep Multimodel SSL DeCUR [102] $\checkmark$ $\checkmark$ $\checkmark$
2023 Oct MSFE+MMFH Feng et al. [27] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Oct ViT FG-MAE [108] $\checkmark$ $\checkmark$ $\checkmark$
2023 Nov ViT Prithvi [46] $\checkmark$ $\checkmark$
2023 Nov Multimodal Encoder CROMA [28] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2023 Dec ViT USat [44] $\checkmark$ $\checkmark$
2024 Jan ViT-B Cross-Scale MAE [88] $\checkmark$ $\checkmark$ $\checkmark$
@@
EarthPT [82] $\checkmark$ $\checkmark$
2024 Jan Teacher-Student Network GeRSP [42] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Jan Dual-Branch SwiMDiff [91] $\checkmark$
2024 Jan Generative ConvNet SMLFR [22] $\checkmark$ $\checkmark$ $\checkmark$
2024 Feb 3D GPT SpectralGPT [40] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Feb MAE-based Framework Presto [92] $\checkmark$ $\checkmark$ $\checkmark$
2024 Mar SatMAE SatMAE++ [73] $\checkmark$ $\checkmark$
@@
SAR-JEPA [58] $\checkmark$ $\checkmark$
2024 Mar ViT FoMo-Bench [8] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
SkySense [32] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Mar Multi-Modules UPetu [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Apr Swim Transformer msGFM [33] $\checkmark$ $\checkmark$ $\checkmark$
2024 Apr DINO DINO-MC [111] $\checkmark$ $\checkmark$ $\checkmark$
2024 May OFA-Net OFA-Net [118] $\checkmark$ $\checkmark$ $\checkmark$
@@
2024 May ViT BFM [11] $\checkmark$ $\checkmark$ $\checkmark$
2024 May MP-MAE MMEarth [72] $\checkmark$ $\checkmark$ $\checkmark$
2024 May ViT CtxMIM [123] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 May HiViT SARATR-X [54] $\checkmark$ $\checkmark$ $\checkmark$
2024 May Transformer SoftCon [106] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 May ViT LeMeViT [49] $\checkmark$ $\checkmark$ $\checkmark$
2024 Jun Masked Autoencoder S2MAE [59] $\checkmark$ $\checkmark$ $\checkmark$
2024 Jun CNN - Transformer RS-DFM [110] $\checkmark$ $\checkmark$
2024 Jun MAE-based A2-MAE [122] $\checkmark$ $\checkmark$ $\checkmark$
2024 Jun ViT HyperSIGMA [95] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
2024 Jun Dynamic OFA DOFA [117] $\checkmark$ $\checkmark$ $\checkmark$
TABLE I: Summary of the pretraining methods utilized and image analysis tasks evaluated across different models. Image-level, pixel-level, region-level, and spatial-temporal classify the tasks in image analysis, while contrastive learning and predictive coding indicates the different self-supervised pretraining strategies that each study used.
In recent years, remote sensing has found applications in many fields. With regard to environmental monitoring, it is used to track deforestation, to monitor air and water quality, and to assess the impacts of climate change [39, 30]. In agriculture, remote sensing helps in crop health monitoring, yield estimation, and efficient resource management [71]. Urban planning and development benefit from remote sensing through the monitoring of urban sprawl, infrastructure development, and land-use planning [48, 17]. Furthermore, in disaster management, remote sensing is crucial for assessing the damage caused by natural disasters, aiding in the planning and execution of relief operations [1, 30].
@@ II-B Foundation Models for Remote Sensing
A major strength of these models lies in their ability to utilize SSL to learn effective representations from largely unlabeled data, which is often abundant in remote sensing scenarios [126]. By integrating advanced architectures like transformers [93], FMs in remote sensing can handle the unique characteristics of geospatial data, such as varying spatial resolutions and temporal dynamics, without requiring separate task-specific models.
The evolution of FMs has been driven by advancements in deep learning and the availability of large datasets. Initially, convolutional neural networks (CNNs) like ResNet [37] paved the way for improved image recognition and classification tasks [65]. The introduction of transformers, which use self-attention mechanisms to model long-range dependencies, has further advanced the capabilities of FMs in handling large-scale image data [16]. Vision transformers (ViTs) [25] extend the transformer architecture to process image data by treating image patches as sequences of tokens, enabling models to learn both local and global relationships. This capability makes transformers particularly effective for semantic segmentation and change detection tasks, where capturing long-range dependencies is crucial, especially in high-resolution satellite imagery.
@@
III Related Review Papers
Artificial intelligence in remote sensing has been a growing area of research, with numerous review papers providing insights into AI advancements and their applications. In this section, we summarize the most influential reviews on foundation models in remote sensing.
Zhang et al. (2016), in their foundational review "Deep Learning for Remote Sensing Data: A Technical Tutorial on the State of the Art" [121], introduced deep learning techniques to RS, focusing on convolutional neural networks (CNNs) for tasks such as image classification and object detection. This work highlighted both the promise and challenges of early AI integration in RS, setting the stage for subsequent advancements.
In 2017, Zhu et al.’s "Deep Learning in Remote Sensing: A Comprehensive Review and List of Resources" [129] delved into diverse AI applications, including hyperspectral analysis and synthetic aperture radar (SAR) interpretation. It also provided an extensive resource list, capturing the rapid adoption of deep learning in addressing complex RS challenges, paving the way for more advanced AI models in the following years.
More recent reviews have focused on advanced AI models and methods. Wang et al.’s 2022 review, "Self-Supervised Learning in Remote Sensing" [103], highlighted the ability of self-supervised learning (SSL) methods to utilize large volumes of unlabeled data, significantly reducing dependence on labeled datasets while maintaining high performance in RS tasks. The review also identified key challenges and future directions, emphasizing SSL’s potential to handle large-scale RS data complexities.
Zhang et al. (2022), in "Artificial Intelligence for Remote Sensing Data Analysis: A Review of Challenges and Opportunities" [120], offered a comprehensive overview of AI algorithms, synthesizing findings from over 270 studies. It emphasized ongoing challenges such as explainability, security, and integrating AI with other computational techniques, serving as a roadmap for future innovation in AI-driven RS.
Aleissaee et al.’s 2023 survey, "Transformers in Remote Sensing" [3], explored the impact of transformer-based models across various RS tasks, comparing them with CNNs. It identified both strengths and limitations, along with unresolved challenges, providing a detailed roadmap for future research on transformers’ role in RS.
Li et al.’s 2024 review, "Vision-Language Models in Remote Sensing" [60], examined the increasing significance of vision-language models (VLMs), which combine visual and textual data. It highlighted VLMs’ potential in applications like image captioning and visual question answering, emphasizing a shift toward richer semantic understanding in RS tasks.
Additionally, the recent work, "On the Foundations of Earth and Climate Foundation Models" [130], provided a comprehensive review of existing foundation models, proposing features like geolocation embedding and multisensory capability. It outlined key traits for future Earth and climate models, contributing to a broader discussion on foundational advancements in geospatial AI.
Building on these reviews, our study provides a comprehensive analysis of foundation models developed from June 2021 to June 2024, focusing on advances in self-supervised learning and transformer-based architectures. Unlike previous reviews, which focused mainly on individual techniques, we explore their combined potential in remote sensing tasks like semantic segmentation, multi-spectral analysis, and change detection. For instance, SatMAE [16] demonstrates effective use of SSL for pre-training transformers, enabling improved segmentation in complex multi-spectral imagery, while Scale-MAE employs scale-aware masked autoencoders for better handling of varied spatial resolutions in remote sensing data.
Our study also highlights new models like DINO-MC [111], which integrates global-local view alignment for SSL, making it particularly effective for identifying changes in high-resolution satellite imagery. By systematically examining these innovations, we illustrate how recent models address persistent challenges like domain adaptation and computational efficiency. For example, efficient self-attention mechanisms in Scale-MAE[78] help reduce computation costs, while enhanced geolocation embeddings in models like SatMAE improve performance in geospatial feature extraction.
In contrast to earlier reviews, which often remained theoretical, we emphasize both the theoretical advancements and practical applications of recent models. For example, DINO-MC [111] and ORBIT’s [101] real-world applications in environmental monitoring and disaster response highlight its practical impact, demonstrating how new FMs can be effectively leveraged to address pressing challenges in geospatial analysis.
IV Pretraining Methods